OpenStudy (anonymous):

MEDAL GIVEN!!!! Rationalize the denominator of sqrt 5 / 6+sqrt2

OpenStudy (solomonzelman):

multiply times the conjugate.

OpenStudy (anonymous):

\[\frac{ \sqrt{5} }{ 6+\sqrt{2} }\]

OpenStudy (anonymous):

to rationalize the denominator, you always multiply by the conjugate

OpenStudy (anonymous):

What is the conjugate

OpenStudy (anonymous):

\[\frac{ \sqrt{5} }{ 6+\sqrt{2} }*\frac{ 6-\sqrt{2}}{ 6-\sqrt{2} }\]

OpenStudy (anonymous):

So then we minus

OpenStudy (anonymous):

the conjugate is when you take the term you have and change the sign sign in the middle as the simplest explanation I can give you. here we have\[6+\sqrt{2}\] so the conjugate is\[6-\sqrt{2}\]

OpenStudy (anonymous):

if we had \[8-\sqrt{13}\] the conjugate would be \[8+\sqrt{13}\]

OpenStudy (anonymous):

Okay so then what do I do

OpenStudy (anonymous):

you multiply the top and bottom by the conjugate

OpenStudy (anonymous):

use the foil method to get your answer

OpenStudy (anonymous):

4.5

OpenStudy (anonymous):

are you asked to give an exact answer or just to rationalize and then leave it?

OpenStudy (anonymous):

Simplify the answer

OpenStudy (anonymous):

A. -2sqrt5 - sqrt30 / 2 B. 6sqrt5-sqrt10/34 C. 6sqrt5-sqrt10/31 D. 5sqrt5 - sqrt12 /23

OpenStudy (anonymous):

well lets multiply it out\[\frac{ 6\sqrt{5}-\sqrt{10} }{ 36-6\sqrt{2}+6\sqrt{2}-\sqrt{2}^{2} }\]

OpenStudy (anonymous):

we cant simplify the top anymore and the 6sqrt2 on the bottom cancel

OpenStudy (anonymous):

Okay would b be the closet answer

OpenStudy (anonymous):

sqrt2^2 equal 2 so we have\[\frac{ 6\sqrt{5}-\sqrt{10} }{ 36-2 }\]\[\frac{ 6\sqrt{5}-\sqrt{10} }{ 34 }\]

OpenStudy (anonymous):

Thank you!!!!

OpenStudy (anonymous):

plugging that into our handy dandy calculator we get a 0.302

OpenStudy (anonymous):

you are welcome